# Verifix: Post-Training Correction to Improve Label Noise Robustness with Verified Samples

### v1

@@ Acknowledgement
This work was supported in part by, the Center for Brain-inspired Computing (C-BRIC), the Center for the Co-Design of Cognitive Systems (COCOSYS), a DARPA-sponsored JUMP center, the Semiconductor Research Corporation (SRC), the National Science Foundation, and DARPA ShELL.
Appendix 0.A Representation Matrix for convolution layer
@@
Appendix 0.B Training and Hyperparameter Details
0.B.1 Sythetic Noise with CIFAR Dataset
For our experiments on CIFAR10 and CIFAR100, we use modified versions of ResNet18 and VGG11 with BatchNorm. All the hyperparameters are kept identical for both datasets. Below we present the hyperparameter details for several baselines.
0.B.1.1 Vanilla Baseline :
@@
0.B.1.2 Finetune Baseline:
We randomly sample 1000 data points from the clean partition of the dataset ($\mathcal{D}_{Tr}^{cln}$) to finetune the Vanilla model using the Post-Training Correction paradigm. For this fine-tuning baseline, we use SGD algorithm for 50 epochs and tune the learning rate using three values [0.005, 0.001, 0.0005]. We also tried using our curvature-based sampling technique for finetuning, however, we found choosing low curvature samples catastrophically decreases the generalization performance of the model.
0.B.1.3 SAM:
All the hyperparameters for the Vanilla baseline are kept the same. We set the $\rho$ parameter to $0.05$ in our experiments.
0.B.1.4 Verifix:
We introduce the scaling hyperparameter $\alpha$ (Equation 8) and perform a search within the range [ 10000, 30000, 100000].
0.B.2 Real World Noise
For our experiments on Mini-WebVision and WebVision1.0, we use InceptionResNetV2. Experiments on Clothing1M are presented on ResNet50 architecture. Below we present the hyperparameter details for several baselines.
0.B.2.1 Vanilla :
We train our models using Stochastic Gradient Descent (SGD) with an initial learning rate of $0.01$. We employ a weight decay of $4\times 10^{-5}$ with Nesterov accelerated momentum with the value of $0.9$. Our experiments on these datasets use a batch size of 64 for Mini-WebVision, 512 for WebVision1.0, and Clothing1M. We randomly initialize the network and train the networks for 200 and 60 epochs for Mini-WebVision and WebVision1.0, respectively. For Clothing1M we initialized the network with Pytorch ResNet50 model pretrained on ImageNet1k. We keep these hyperparameters the same for the rest of the baselines.
0.B.2.2 MixUp :
@@ 0.B.2.4 SAM :
0.B.2.5 Verifix :
We introduce the scaling hyperparameter $\alpha$ (Equation 8) and perform a search within the range [ 10000, 30000, 100000].
Appendix 0.C Experiments with low-level synthetic noise
We conduct experiments with varying degrees of label corruption on the CIFAR10 and CIFAR100 datasets, as presented in Table 3.
Table 3: Test Accuracy for label noise correction on CIFAR10 and CIFAR100 datasets trained on VGG11 and ResNet18 architectures with varying amounts of label corruption. Finetuning and Our method assumes access to 1000 trusted or expert-verified samples.
Method VGG11_BN ResNet18
Percentage Mislabeled Samples Percentage Mislabeled Samples
2% 5% 10% 25% 2% 5% 10% 25%
CIFAR10 Ideal $91.18\pm 0.17$ $94.23\pm 0.06$
Retrain $90.93\pm 0.12$ $91.0\pm 0.21$ $90.73\pm 0.15$ $89.6\pm 0.21$ $93.96\pm 0.02$ $94.07\pm 0.12$ $93.77\pm 0.25$ $93.08\pm 0.21$
Vanilla $89.89\pm 0.12$ $88.16\pm 0.14$ $85.01\pm 0.38$ $72.19\pm 0.33$ $93.13\pm 0.15$ $91.54\pm 0.17$ $88.66\pm 0.38$ $78.37\pm 0.18$
Finetune $90.07\pm 0.09$ $88.79\pm 0.20$ $87.03\pm 0.42$ $81.82\pm 0.25$ $93.23\pm 0.10$ $91.84\pm 0.18$ $89.49\pm 0.2$ $82.82\pm 0.70$
Verifix $\textbf{90.11}\pm\textbf{0.23}$ $\textbf{89.21}\pm\textbf{0.04}$ $\textbf{87.78}\pm\textbf{0.18}$ $\textbf{83.38}\pm\textbf{0.47}$ $\textbf{93.26}\pm\textbf{0.13}$ $\textbf{92.16}\pm\textbf{0.15}$ $\textbf{90.46}\pm\textbf{0.10}$ $\textbf{86.96}\pm\textbf{0.14}$
CIFAR100 Ideal $67.44\pm 0.18$ $75.34\pm 0.06$
Retrain $66.85\pm 0.49$ $67.05\pm 0.13$ $65.99\pm 0.43$ $63.73\pm 0.52$ $74.69\pm 0.4$ $74.65\pm 0.29$ $73.89\pm 0.16$ $72.13\pm 0.36$
Vanilla $66.02\pm 0.5$ $63.75\pm 0.55$ $60.66\pm 0.23$ $49.03\pm 0.69$ $74.19\pm 0.33$ $71.79\pm 0.46$ $68.16\pm 0.40$ $57.60\pm 0.17$
Finetune $66.02\pm 0.50$ $63.76\pm 0.56$ $60.68\pm 0.23$ $49.01\pm 0.70$ $\textbf{74.20}\pm\textbf{0.23}$ $71.86\pm 0.50$ $68.42\pm 0.53$ $58.67\pm 0.17$
Verifix $\textbf{66.57}\pm\textbf{0.40}$ $\textbf{65.09}\pm\textbf{0.56}$ $\textbf{63.10}\pm\textbf{0.40}$ $\textbf{56.81}\pm\textbf{0.37}$ $74.18\pm 0.31$ $\textbf{72.03}\pm\textbf{0.42}$ $\textbf{69.16}\pm\textbf{0.16}$ $\textbf{61.42}\pm\textbf{0.46}$

### v2

@@ Acknowledgements
This work was supported in part by the Center for the Co-Design of Cognitive Systems (COCOSYS), a DARPA-sponsored JUMP center, the Semiconductor Research Corporation (SRC), the National Science Foundation, and Collins Aerospace.
Appendix A Experimental Setup for Toy Illustration
In Figure 2 we demonstrate our unlearning algorithm on a 2 way classification problem, where the original model is trained to detect samples from 2 different 2-dimensional spirals marked by green and blue points in Figure 2(a). Below we present the details of the experiment.
Dataset:
The training dataset has 250 samples per class and the test dataset has 5000 samples per class.
Model Architecture:
We use a simple linear model with 10 layers with with ReLU activation functions. All the intermediate layers have 500 neurons and each layer excluding the final layer is followed by BatchNorm.
Training details:
We train this network with stochastic gradient descent for 250 epochs with a learning rate of 0.01. We use learning rate scheduler which scales the learning rate by 0.7 when the training loss has stopped reducing. For these experiment we use a training batch size of 512.
Appendix B Brief Literature Survey on Label Noise Learning
(a) Dataset cleaning
(b) Noise-Robust Training
(c) Corrective Unlearning
Figure 7: Overview of Label Noise Handling. (a) Explores Data Cleaning approaches, (b) Presents Noise-Robust Training methods and (c) shows Corrective Unlearning.
Data Cleaning: One prominent direction to handle label noise involves data filtering, where the goal is to identify and remove mislabeled samples within the training data. Ensemble learning approaches (Feng, Quan, and Dauphin 2020) have been explored for filtering. However, these methods can be computationally expensive due to the need to train multiple models. To address this challenge, recent studies have focused on leveraging training dynamics to identify label noise, including methods using learning time (Xia et al. 2020), forgetting time (Maini et al. 2022), train-set confidence (Northcutt, Jiang, and Chuang 2021) or input loss curvature (Garg, Ravikumar, and Roy 2024; Ravikumar et al. 2024). These offline filtering often require the retraining of the network on the clean dataset. Researchers have creatively solved the above challenge with sample selection (Cheng et al. 2021; Patel and Sastry 2023) techniques, where the algorithm dynamically removes samples from the training data. Since both filtering and sample selection can reduce training data size, potentially impacting model performance, research has also explored label correction (Zheng et al. 2020; Zheng, Awadallah, and Dumais 2021) algorithms.
Noise-Robust Training: Several works focus on developing algorithms that are inherently less sensitive to label noise. Regularization techniques, such as dropout (Jindal, Nokleby, and Chen 2016), label smoothing (Lukasik et al. 2020; Wei et al. 2021) or clipping the model predictions (Wei et al. 2023), help prevent overfitting to noisy labels, leading to improved robustness. Noise robust loss functions like symmetric cross-entropy (Wang et al. 2019) and Mean Absolute Error (MAE) less sensitive to outliers caused by noisy labels compared to standard cross-entropy loss. Further, Optimization methods like Sharpness-Aware Minimization (SAM) (Foret et al. 2021) promote flatter loss landscapes, which are associated with better generalization under label noise. Additionally, curriculum learning (Jiang et al. 2018; Zhang and Sabuncu 2018) gradually increase the difficulty of training examples, helping the model learn robust features before encountering noise. Data augmentation techniques such as MixUp (Zhang et al. 2017), DivideMix (Li, Socher, and Hoi 2020), MentorMix(Jiang et al. 2020) smooth training distributions, encouraging less abrupt decision boundaries and further enhancing robustness. Further, several studies estimate the Noise Transition Matrix (Zhu, Wang, and Liu 2022; Cheng et al. 2022) to incorporate the label noise into the training process. While these robust algorithms offer improvement over standard training, however, do not completely close the generalization gap compared to the model trained on clean data.
Appendix C Representation Matrix for convolution layer
Figure 8 illustrates the unfold operations in details.
@@ Appendix C Representation Matrix for convolution layer
Appendix D Training and Hyperparameter Details
Sythetic Noise with CIFAR Dataset
For our experiments on CIFAR10 and CIFAR100, we use modified versions of ResNet18 and VGG11 with BatchNorm. All the hyperparameters ( or hyperparameter tuning ranges ) are kept identical for both datasets. Below we present the hyperparameter details for several baselines.
Ideal/Retrain/Vanilla Baseline :
@@ Appendix D Training and Hyperparameter Details
Finetune Baseline:
We randomly sample 5000 data points from the clean partition of the dataset ($\mathcal{D}_{Tr}^{cln}$) to finetune the Vanilla model using the Post-Training Correction paradigm. For this fine-tuning baseline, we use SGD algorithm for 50 epochs and tune the learning rate using three values [0.005, 0.001, 0.0005]. We also tried using our loss-based sampling technique for finetuning as our method, however, we found choosing low loss samples catastrophically decreases the generalization performance of the model.
SSD:
SSD has the hyperparameter $\alpha$ which is tuned in the ranges [0.1, 1, 10, 50, 100, 500, 1000, 1e4, 1e5, 1e6] and hyperparameter $\lambda=\beta\times\alpha$ where $\beta$ is in the range [0.1, 0.5, 1.0, 5, 10] following (Goel et al. 2024).
SCRUB:
We tune the following hyperparameter for SCRUB:
  • retain-batch-size - [32, 64, 128]
  • forget-batch-size - [32, 64, 128]
  • $\alpha$ [0.03, 1, 0.3, 1, 3]
  • $\gamma$ [0.03]
  • number of minimization steps - [5, 10, 15, 20, 25]
  • number of maximization steps - [1, 3, 10]
SAP:
We introduce the scaling hyperparameter $\alpha$ (Equation 7) and perform a search within the range [2000, 4000, 8000, 10000, 12500, 15000, 17500, 20000, 22500, 25000, 30000, 40000, 50000, 75000, 100000, 300000].
MixUp :
@@
Logit Clip:
We tune the hyperparameter temperature for logit clipping in [0.01, 0.02, 0.04, 0.06, 0.08, 1.0, 1.5, 2, 2.5, 3.0, 3.5, 4.0, 4.5, 5.0] and found $3.0$ gave better results.
Real World Noise
For our experiments on Mini-WebVision we use InceptionResNetV2. Experiments on Clothing1M are presented on ResNet50 and ViT_B_16 architecture. Below we present the hyperparameter details for several baselines.
Vanilla :
For convolutional networks, we train our models using Stochastic Gradient Descent (SGD) with an initial learning rate of $0.01$. We employ a weight decay of $4\times 10^{-5}$ with Nesterov accelerated momentum with the value of $0.9$. Our experiments on these datasets use a batch size of 64 for Mini-WebVision, and Clothing1M. We randomly initialize the network and train the networks for 200 and 60 epochs for Mini-WebVision and WebVision1.0, respectively. For Clothing1M we initialized the network with Pytorch ResNet50 model pretrained on ImageNet1k. We keep these hyperparameters the same for the rest of the baselines.
For the transformer based model (ViT_B_16), we use the learning rate of $2^{-5}$ and weight decay of 0.01 with a nestrove momentum of 0.9. We initialize the network with the pretrained ImageNet1k weights. We use the batchsize of 1024 and train the model for 10 epochs.
SAP :
We tune the scaling hyperparameter $\alpha$ (Equation 7) within the range [$5e^{3}$, $1e^{4}$, $2e^{4}$, $3e^{4}$, $4e^{4}$, $6e^{4}$, $8e^{4}$, $1e^{5}$, $2e^{5}$, $4e^{5}$, $6e^{5}$, $8e^{5}$, $1e^{6}$, $2e^{6}$, $4e^{6}$, $6e^{6}$ ].
Appendix E Additional Results for SCRUB
SAP outperforms all unlearning methods on both CIFAR10 and CIFAR100 datasets with and average improvement of $1.36\%$ and $0.39\%$ in comparison to the second best method as presented in Table 1. We observe that increasing the number of retain samples to 5000 and forget samples to 1000 increases the performance of SCRUB algorithm making SCRUB competitive with SAP for CIFAR10 and marginally ($0.71\%$) superior to SAP for CIFAR100 dataset. However, SCRUB requires extensive hyperparameter tuning (retain batch size, forget batch size, $\alpha$, $\gamma$, number of minimization steps, and number of maximization steps) to counteract the instability of gradient ascent, leading to approximately 675 hyperparameter search iterations compared to our method’s 16. Increasing the number of samples further exacerbates the computational burden for hyperparameter search.
  Method VGG11_BN ResNet18 Average
Retain Forget Symmetric Noise ( Eqn. 9 ) Symmetric Noise ( Eqn. 9 ) Asymmetric Noise ( Eqn. 10 ) Hierarchical Noise ( Eqn. 11 )
samples samples $\eta=0.1$ $\eta=0.25$ $\eta=0.1$ $\eta=0.25$ $\eta=0.1$ $\eta=0.25$ $\eta=0.1$ $\eta=0.25$
CIFAR10 SCRUB $1000$ $200$ $85.88\pm 0.35$ $78.90\pm 0.25$ $89.50\pm 0.17$ $83.77\pm 0.44$ $89.50\pm 0.22$ $83.60\pm 0.14$ $91.65\pm 0.12$ $88.00\pm 0.58$ $86.35$
SCRUB $5000$ $1000$ $87.05\pm 0.13$ $81.60\pm 0.39$ $89.99\pm 0.17$ $85.33\pm 0.30$ $\textbf{90.19}\pm\textbf{0.18}$ $85.68\pm 0.78$ $\textbf{92.06}\pm\textbf{0.38}$ $\textbf{89.29}\pm\textbf{0.34}$ $87.65$
SAP 00 00 $\textbf{87.25}\pm\textbf{0.16}$ $\textbf{82.27}\pm\textbf{0.15}$ $\textbf{90.12}\pm\textbf{0.11}$ $\textbf{85.49}\pm\textbf{0.39}$ $90.03\pm 0.25$ $\textbf{86.32}\pm\textbf{0.66}$ $91.87\pm 0.22$ $87.92\pm 0.69$ 87.66
CIFAR100 SCRUB $1000$ $200$ $60.93\pm 0.09$ $52.11\pm 0.63$ $67.02\pm 0.29$ $57.36\pm 0.40$ $73.12\pm 0.18$ $63.37\pm 0.69$ $68.20\pm 0.13$ $60.24\pm 0.33$ $62.79$
SCRUB $5000$ $1000$ $\textbf{61.50}\pm\textbf{0.23}$ $\textbf{54.54}\pm\textbf{0.96}$ $\textbf{67.60}\pm\textbf{0.26}$ $\textbf{59.23}\pm\textbf{0.44}$ $\textbf{73.24}\pm\textbf{0.53}$ $\textbf{64.10}\pm\textbf{0.64}$ $\textbf{68.83}\pm\textbf{0.28}$ $\textbf{62.10}\pm\textbf{0.06}$ 63.89
SAP 00 00 $61.10\pm 0.23$ $53.31\pm 0.78$ $66.82\pm 0.17$ $58.74\pm 0.61$ $72.92\pm 0.30$ $63.57\pm 0.49$ $68.24\pm 0.17$ $60.76\pm 0.50$ $63.18$
Table 4: Test Accuracy for symmetric noise (Equation 9) removal from CIFAR10 and CIFAR100 datasets trained on VGG11 and symmetric noise, asymmetric noise (Equation 10) and hierarchical noise (Equation 11) removal ResNet18 architectures with noise strength $\eta=0.1$ and $\eta=0.25$. Note, the confusing class groups (or set $\mathcal{C}$) for hierarchical noise are cat-dog and truck-automobile for CIFAR10. For CIFAR100 we use Superclasses as the confusing groups.
- We select 1000 samples with low loss value from $\mathcal{D}_{Tr}$ as retain samples for our method.